Find the area bounded by the parabola y2 = 4ax and its latus rectum.


1
Expert's answer
2021-05-07T12:29:25-0400



Given parabola is "y^2 = 4ax"


 Its focus is "(a, 0)"


 Equation of latus rectum is "x = a"


The parabola is symmetrical about the x-axis  


Required area = "2[\text{ Area in the first quadrant between the limits } x = 0 \text{ and } x = a]"



"=2\int_0^aydx\\[9pt]=2\int_0^a\sqrt{4ax}dx\\[9pt]=2(2\sqrt{a})\int_0^ax^{\frac{1}{2}}dx



\\[9pt]=4\sqrt{a}[\dfrac{x^{\frac{3}{2}}}{\frac{3}{2}}]|_0^a



\\[9pt]=4\sqrt{a}(\dfrac{8}{3}a^2)=\dfrac{8}{3}a^2 sq. units"